MarcoValtorta 1.PersonalHistory OFFICEADDRESS SumwaltRoom329 DepartmentofComputerScienceandEngineering UniversityofSouthCarolina Columbia,SouthCarolina29208 U.S.A. tel.(1)(803)777-4641 HOMEADDRESS 4729DaturaRoad Columbia,SouthCarolina29205 U.S.A. tel.(1)(803)782-9898 BIOGRAPHICALDATA BornonMay7,1956,Milan,Italy Italiancitizen,permanentresidentoftheU.S.sinceAugust1988 MarriedtoaU.S.citizensince1982.Onedaughterandoneson. RESEARCHINTERESTS Maininterests:ArtificialIntelligence,especiallyprobabilisticreasoning,expertsystems,problemrepresentationandsearch,heuristics.Otherinterests:complexitytheory,designofefficient programs,logicprogramming,VLSIalgorithmics. EDUCATION Ph.D.,ComputerScience,DukeUniversity,April1987. Dissertation:"AutomatingRuleStrengthsinExpertSystems." Supervisor:D.W.Loveland. M.A.,ComputerScience,DukeUniversity,April1983. Project:"TheGraduateCourseAdviser." Supervisor:D.W.Loveland. LaureainIngegneriaElettronica(cumlaude--highesthonorawarded),PolitecnicodiMilano,July1980. Thesis:"UncontributoallaTeoriadellaRisoluzionedeiProblemi:RappresentazioneSemantica,ProprietaAlgebricheeAlgoritmidiRicerca." Supervisor:MarcoSomalvico. LANGUAGES Italian(firstlanguage),English(veryfluent),French,German(some),Latin,AtticGreekof the4thcenturyB.C.(studied). 2.Non-AcademicExperience October1985-August1988 ProjectOfficerinESPRIT,aprogramoftheEuropeanEconomicCommunity.Dutiesincluded supervisionofsponsoredprojects,negotiationofshortlistedproposals,participationinthe evaluationofproposalsandinthedevelopmentoftheyearlyworkprogram.Isupervisedan averageofnineprojectsinthefieldofknowledgeengineering,withanaverageof4to5participatingorganizationsperproject.Iindependentlynegotiatedfiveproposalsintoprojectsand participatedinthenegotiationprocessforseveralmore. 3.AcademicProfessionalExperience June1994-present AssociateProfessorattheDepartmentofComputerScienceandEngineering1,Universityof SouthCarolina. June1993-August1999 DirectorofUndergraduateStudiesattheDepartmentofComputerScience,UniversityofSouth Carolina. August1988-June1994 AssistantProfessorattheDepartmentofComputerScience,UniversityofSouthCarolina. August1980-August1985 InstructionalandResearchAssistantforProfessorD.W.LovelandatDukeUniversity. 4.PublicationsinRefereedJournals Valtorta,Marco,JiriVomlel,andYoung-GyunKim."SoftEvidentialUpdateforMultiagent Systems."InternationalJournalofApproximateReasoning.Underreview. Mani,Subramani,SuzanneMcDermott,andMarcoValtorta."MENTOR:ABayesianModel forPredictionofMentalRetardationinNewborns."ResearchinDevelopmentalDisabilities, 18(1997),5,pp.303-318. Ling,C.X.F.andM.Valtorta."RefinementofUncertainRuleBasesviaReduction."InternationalJournalofApproximateReasoning,13,2(August1995).95-126. Singh,M.andM.Valtorta."ConstructionofBayesianBeliefNetworksfromData:aBriefSurveyandanEfficientAlgorithm."InternationalJournalofApproximateReasoning,12,2 (February1995),111-131. Valtorta,M.andM.I.Zahid."Tie-BreakingRulesfor4xnWarnsdorff’sTours."Congressus Numerantium,95(1993),75-86. Valtorta,M.andM.I.Zahid."Warnsdorff’sToursofaKnight."JournalofRecreational Mathematics,25(1993),2,128-140. Valtorta,M.andD.W.Loveland."OntheComplexityofBeliefNetworkSynthesisandRefinement."InternationalJournalofApproximateReasoning,7,3-4(October-November1992), 121-148. 1DepartmentofComputerScienceuntil2000.

Hansson,O.,A.Mayer,andM.Valtorta."ANewResultontheComplexityofHeuristicEstimatesfortheA*Algorithm."ArtificialIntelligence,55,1(May1992),129-143. Durham,S.D.,J.S.Smolka,andM.Valtorta."StatisticalConsistencywithDempster’sRule onDiagnosticTreesHavingUncertainPerformanceParameters."InternationalJournalof ApproximateReasoning,6,1(January1992),67-81. Valtorta,M."Knowledge-BasedConstructionofProbabilisticandDecisionModels"(AAAI-91 workshopreport).IEEEExpert,6,6(December1991),69-71. Valtorta,M."KnowledgeBaseRefinement:ABibliography."AppliedIntelligence,1,1(July 1991),87-94. Valtorta,M."SomeResultsontheComputationalComplexityofRefiningConfidenceFactors."InternationalJournalofApproximateReasoning,5,2(March1991),123-148. Virtanen,H.,K.Johnson,R.A.O’Keefe,andM.Valtorta."FuzzyProlog."Newsletterofthe AssociationforLogicProgramming,4,1(February1991),11-12. Valtorta,M."AResultontheComputationalComplexityofHeuristicEstimatesfortheA* Algorithm."InformationSciences,vol34(1984),47-59. 5.InvitedSurveys/Tutorials/ PapersPublishedinConferenceProceedings/Books PublishedBookChapter Loveland,D.W.andM.Valtorta."DetectingAmbiguity:AnExampleinKnowledgeEvaluation."In:Gupta,U.(ed.).ValidatingandVerifyingKnowledge-BasedSystems.LosAlamitos, CA:IEEEComputerSocietyPress,391-395,1991. 6.ContributedSurveys/Tutorials/ PapersPublishedinConferenceProceedings2/Books Ghosh,JayantaK.andM.Valtorta."ProbabilisticBayesianNetworkModelBuildingofHeart Disease(ExtendedAbstract)."Proceedingsofthe38thAnnualACMSoutheasternConference, Clemson,SouthCarolina,April7-8,2000. Bloemeke,MarkandM.Valtorta."AHybridAlgorithmtoComputeMarginalandJoint BeliefsinBayesianNetworksandItsComplexity."In:G.F.CooperandS.Moral(eds.), UncertaintyinArtificialIntelligence:ProceedingsoftheFourteenthConference.SanFrancisco, CA:Morgan-Kaufmann,1998,16-23.(*) Kim,Young-GyunandM.Valtorta."OntheDetectionofConflictsinDiagnosticBayesian 2Someofthepapersarepublishedin"workingnotes,"asindicated.Workingnotesareoftenavailableonlytoparticipantsofa workshoporconference.Refereedpaperspublishedinproceedingsaremarkedwithanasterisk(*).Notethatsomeofthepapers publishedinworkingnotesarerefereed,eventhoughnonearemarkedwithanasterisk.

NetworksUsingAbstraction."In:Ph.BesnardandS.Hanks(eds.),UncertaintyinArtificial Intelligence:ProceedingsoftheEleventhConference.SanFrancisco,CA:Morgan-Kaufmann, 1995,362-367.(*) Subramani,Mani,M.Valtorta,andS.McDermott."MENTOR:ABayesianModelforPredictionandInterventioninMentalRetardation."WorkingNotesoftheFifthInternationalWorkshoponArtificialIntelligenceandStatistics(AIS-95),366-371,Ft.Lauderdale,FL,January 1995. Mechling,RandyandM.Valtorta."AParallelConstructorofMarkovNetworks."Pages 255-262in:Cheeseman,PeterandWayneOldford(eds.):SelectingModelsfromData:Artificial IntelligenceandStatisticsIV.Volume89ofLectureNotesinStatistics.NewYork:Springer-Verlag,1994(ISBN0-387-94281-5). Bordetski,AlexanderandM.Valtorta."LearningEmpiricalConstraintstoComplementDiagnosticModels."In:C.H.Dagli,L.I.Burke,B.R.Fernandez,J.Ghosh(eds.),IntelligentEngineeringSystemsthroughArtificialNeuralNetworks,Vol.3(ProceedingsoftheThirdInternationalConferenceonArtificialNeuralNetworksinEngineering(ANNIE-93)),97-102,New York,NY:ASMEPress,1993.(*) Childress,RitaL.andM.Valtorta."Polynomial-TimeModel-BasedDiagnosiswiththeCriti-calSetAlgorithm."WorkingNotesoftheFourthInternationalWorkshoponPrinciplesof Diagnosis(DX-93),166-177,Aberystwyth,Wales,September1993. Singh,MoninderandM.Valtorta."AnAlgorithmfortheConstructionofBayesianNetwork StructuresfromData."In:D.HeckermanandE.H.Mamdani(eds.),UncertaintyinArtificial Intelligence:ProceedingsoftheEighthConference.SanMateo,CA:Morgan-Kauffman,1993, 259-265.(*) Mechling,Randy,andMarcoValtorta."PaCCIN:AParallelConstructorofMarkovNetworks."WorkingNotesoftheFourthInternationalWorkshoponArtificialIntelligenceand Statistics(AIS-93),405-410,Ft.Lauderdale,FL,January1993. Childress,RitaL.,M.Valtorta,andGiorgioTornielli."TheComplexityofDiagnosingContinuousProcesseswithODS-PI."WorkingNotesoftheAAAIWorkshoponApproximationsand AbstractionsofComputationalTheories,39-46,SanJose,CA,July1992. Wang,ShijieandM.Valtorta."OnTheExponentialGrowthRateofDempster-ShaferBelief Functions."ProceedingsoftheSPIEConferenceonApplicationsofArtificialIntelligenceX: Knowledge-BasedSystems,15-24,Orlando,FL,April1992.(*) Wang,S.andM.Valtorta."OntheConversionofRuleBasesintoBeliefNetworks."Proceedingsofthe1992ACM/SIGAPPSymposiumonAppliedComputing,363-368,KansasCity, MO,March1992.(*) Childress,R.L.andM.Valtorta."EVAandtheVerificationofExpertSystemsWrittenin OPS-5."WorkingNotesoftheAAAI-91WorkshoponKnowledge-basedSystemsVerification, Validation,andTesting,Anaheim,CA,July1991,72-83. Valtorta,M."ComplexityofKnowledgeBaseRefinement(ResearchSummary)."Working NotesoftheAAAI-91WorkshoponKnowledge-BasedConstructionofProbabilisticandDeci-sionModels,Anaheim,CA,July1991,145-148.

Ling,Xiaofeng(Charles)andM.Valtorta."RevisionofReducedTheories."InBirnbaum,L. andG.Collins(eds.)MachineLearning:ProceedingsoftheEighthInternationalWorkshop (IWML-91).SanMateo,CA:MorganKaufmann,1991,519-523.(*) Valtorta,M.andM.I.Zahid."SomeHeuristicsCannotbeDerivedfromSimplifiedModels (abstract)."ProceedingsoftheACM1991ComputerScienceConference,SanAntonio,TX, March5-8,1991,p.698. Durham,S.D.,J.S.Smolka,andM.Valtorta."ConditionsfortheStatisticalConsistencyof Dempster’sRule."WorkingnotesoftheThirdInternationalWorkshoponArtificialIntelligenceandStatistics,Ft.Lauderdale,Florida,January1991,7.1-7.4. Valtorta,M.andM.I.Zahid."SomeHeuristicsCannotbeDerivedfromSimplifiedModels." WorkingNotesoftheAAAI-90WorkshopontheAutomaticGenerationofApproximationsand Abstractions,Boston,Massachusetts,July1990. Wang,S.andM.Valtorta."APrototypeBeliefNetwork-basedExpertSystemShell,"Pro-ceedingsoftheThirdInternationalConferenceonIndustrialandEngineeringApplicationsof ArtificialIntelligenceandExpertSystems(IEA/AIE-90),Charleston,SouthCarolina,July 1990,509-518.(*) Valtorta,M."MoreResultsontheComplexityofKnowledgeBaseRefinement:BeliefNetworks,"ProceedingsoftheSeventhInternationalConferenceonMachineLearning(ICML-90), Austin,Texas,June1990,419-426.(*) Childress,R.andM.Valtorta."VerificationandValidationofExpertSystems."Proceedings oftheSixthAnnualUSC-CSSymposium:IntelligentSystems,Columbia,SouthCarolina, March1990,55-68. Valtorta,M.andM.I.Zahid."OnaConjecturebyJudeaPearl:FirstResults."Proceedingsof theSixthAnnualUSC-CSSymposium:IntelligentSystems,Columbia,SouthCarolina,March 1990,45-54. Valtorta,M."KADSvs.KEATS."ProceedingsoftheIJCAI-89WorkshoponKnowledge Acquisition:PracticalToolsandTechniques,August1989. Valtorta,M."SomeResultsontheComplexityofKnowledge-BaseRefinement."Proceedings oftheSixthInternationalWorkshoponMachineLearning(IWML-89),Ithaca,NewYork,June 1989,326-331.(*) Valtorta,M."AutomatingRuleStrengthsinExpertSystems."Proceedingsofthe8thEuropeanConferenceonArtificialIntelligence(ECAI-88),Munich,Germany,August1988, 369-371.(*) Valtorta,M.,B.T.Smith,andD.W.Loveland."TheGraduateCourseAdvisor:AMulti-Phase Rule-BasedExpertSystems."ProceedingsoftheIEEEWorkshoponPrinciplesofKnowledge-Basedsystems,Denver,Colorado,December1984,53-57.(*) Valtorta,M."KnowledgeRefinementinRuleBasesforExpertSystems:AnApplication-DrivenApproach."ProceedingsoftheFirstInternationalWorkshoponExpertDatabaseSys-tems,KiawahIsland,SouthCarolina,October1984. M.Valtorta."AResultontheComputationalComplexityofHeuristicEstimatesfortheA* Algorithm."Proceedingsofthe8thInternationalJointConferenceonArtificialIntelligence (IJCAI-83),Karlsruhe,Germany,August1983,777-779.(*) Loveland,D.W.andM.Valtorta."DetectingAmbiguity:AnExampleinKnowledgeEvaluation."Proceedingsofthe8thInternationalJointConferenceonArtificialIntelligence (IJCAI-83),Karlsruhe,Germany,August1983,182-184.(*) Valtorta,M."AResultontheComputationalComplexityofHeuristicEstimatesfortheA* Algorithm(ExtendedAbstract)."Proceedingsofthe21stSoutheastRegionACMConference, Durham,NorthCarolina,April1983. Guida,G.,M.Somalvico,andM.Valtorta."ProblemiAusiliariedAlgoritmidiRicerca:Un ContributoallaTeoriadeiProblemi."AttidelCongressoAnnualeAICA-80.Bologna,Italy: Tecnoprint,1980.(*) 7.ContractTechnicalReports Valtorta,Marco,Young-GyunKim,andJiriVomlel."SoftEvidentialUpdateforCommunica-tioninMultiagentSystemsandtheBigCliqueAlgorithm."October31,2000. Childress,RitaL.andMarcoValtorta."DeliverableReportD6:FinalReport."Sixthreport ofproject"AStudyofthecomplexityofAbstractioninQualitativeDiagnosis."June30, 1994.Fourpagesanda3.5"disketteintarformat. Childress,RitaL.andMarcoValtorta."DeliverableReportD5:ProgressonNewDiagnostic DiagnosticAlgorithmsandPrograms."Fifthreportofproject"AStudyoftheComplexityof AbstractioninQualitativeDiagnosis."June7,1993,62pagesnotincludingappendices. Childress,RitaL.andMarcoValtorta."DeliverableReportD4:ProgressonNewDiagnostic Problems."Fourthreportofproject"AStudyoftheComplexityofAbstractioninQualitative Diagnosis."November19,1992,48pages. Matthews,M.M.,A.Sengupta,andM.Valtorta."SentencingInformationRetrievalandCriminalOffenseCoding:FinalReport."August1992,55pages. Childress,RitaL.andMarcoValtorta."DeliverableReportD3:NewDiagnosticProblems." Thirdreportofproject"AStudyoftheComplexityofAbstractioninQualitativeDiagnosis." July9,1992,68pagesnotincludingappendices. Childress,RitaL.andMarcoValtorta."DeliverableReportD2:ComplexityofODS-PIand RelatedIssues."Secondreportofproject"AStudyoftheComplexityofAbstractioninQualitativeDiagnosis."October3,1991,41pagesnotincludingappendices. Childress,RitaL.andMarcoValtorta."DeliverableReportD1:SurveyoftheStateoftheArt andProblemFormalizationinQualitativeDiagnosticReasoningandComplexityTheory." Firstreportofproject"AStudyoftheComplexityofAbstractioninQualitativeDiagnosis." May31,1991,88pagesnotincludingappendices. 8.Non-ContractTechnicalReports Valtorta,M.,Young-GyunKim,andJiriVomlel."SoftEvidentialUpdateforProbabilistic Multi-AgentSystems."TechnicalReportTR0001,DepartmentofComputerScienceandEngi-neering,UniversityofSouthCarolina,2000. Ghosh,JayantaandM.Valtorta"ProbabilisticBayesianNetworkModelBuildingofHeartDisease."TechnicalReportTR9911,DepartmentofComputerScience,UniversityofSouthCarolina,1999. Smith,Wayne,MarkBloemeke,andM.Valtorta."MessagesforAgent-EncapsulatedBayesian NetworksandHowtoUpdateBeliefswithThem."TechnicalReportTR9907,Departmentof ComputerScience,UniversityofSouthCarolina,1999. SubramaniMani,M.Valtorta,andSuzanneW.McDermott."MENTOR:AStudyinBuilding BayesianModelsinMedicine."TechnicalReportTR9601,DepartmentofComputerScience, UniversityofSouthCarolina,1996. SubramaniMani,M.Valtorta,andSuzanneW.McDermott."BayesianModelBuildingfroma LargeMedicalDataset:APracticalStudy."TechnicalReportTR9502,DepartmentofComputerScience,UniversityofSouthCarolina,1995. Kim,Young-GyunandM.Valtorta."OntheDetectionofConflictinBayesianNetworks UsingAbstraction."TechnicalReportTR9501,DepartmentofComputerScience,University ofSouthCarolina,1995. Bordetski,AlexanderandM.Valtorta."LearningEmpiricalConstraintstoComplementDiagnosticModels."TechnicalReportTR9303,DepartmentofComputerScience,Universityof SouthCarolina,1993. Childress,RitaL.andM.Valtorta.‘‘Polynomial-TimeModel-BasedDiagnosiswiththeCriti-calSetAlgorithm.’’TechnicalReportTR9302,DepartmentofComputerScience,Universityof SouthCarolina,1993. Singh,MoninderandM.Valtorta."AnAlgorithmfortheConstructionofBayesianNetwork StructuresfromData."TechnicalReportTR9301,DepartmentofComputerScience,UniversityofSouthCarolina,1993. Valtorta,M.andM.I.Zahid."SomeHeuristicsCannotbeDerivedfromRelaxedModels." TechnicalReportTR9204,DepartmentofComputerScience,UniversityofSouthCarolina, 1992.(ThisreportsupercedesTR90007.) Childress,R.L.,M.Valtorta,andG.Tornielli."OntheComplexityofDiagnosingContinuous ProcesseswithODS-PI."TechnicalReportTR9201,DepartmentofComputerScience,Uni-versityofSouthCarolina,1992. Wang,S.andM.Valtorta."OntheExponentialGrowthRateofDempster-ShaferBeliefFunc-tions."TechnicalReportTR9109,DepartmentofComputerScience,UniversityofSouthCar-olina,1991. Childress,R.L.,M.Valtorta,andG.Tornielli."OntheComplexityofDiagnosingContinuous Processes."TechnicalReportTR9107,DepartmentofComputerScience,UniversityofSouth Carolina,1991. Yu,E.K.andM.Valtorta."Computer-assistedKnowledgeEngineering:ASurveyofKADS andKEATS."TechnicalReportTR9104,DepartmentofComputerScience,Universityof SouthCarolina,1991. Hansson,O.,A.Mayer,andM.Valtorta."ANewResultontheComplexityofHeuristicEstimatesfortheA*Algorithm."TechnicalReportTR9102,DepartmentofComputerScience, UniversityofSouthCarolina,1991. Ling,X.C.andM.Valtorta."ATheoryofRuleBaseRefinementviaTheoryReduction." TechnicalReportTR9101,DepartmentofComputerScience,UniversityofSouthCarolina, 1991. Durham,S.D.,J.S.Smolka,andM.Valtorta."ConditionsfortheConsistencyofDempster’s RuleinDiagnosticTrees."TechnicalReportTR90008,DepartmentofComputerScience,UniversityofSouthCarolina,1990. M.ValtortaandM.I.Zahid."SomeHeuristicsCannotbeDerivedfromSimplifiedModels." TechnicalReportTR90007,DepartmentofComputerScience,UniversityofSouthCarolina, 1990. Wang,S.andM.Valtorta."APrototypeBelief-NetworkBasedExpertSystemShell."Tech-nicalReportTR90006,DepartmentofComputerScience,UniversityofSouthCarolina,1990. Childress,R.L.andM.Valtorta."PreliminaryReportontheVerificationofExpertSystems WritteninOPS5."TechnicalReportTR90005,DepartmentofComputerScience,University ofSouthCarolina,1990. M.ValtortaandD.W.Loveland."OntheComplexityofBeliefNetworkSynthesisandRefinement."TechnicalReportTR89011,DepartmentofComputerScience,UniversityofSouth Carolina,1989(updatedJune1991). M.Valtorta."KADSvs.KEATS."TechnicalReportTR89009,DepartmentofComputerScience,UniversityofSouthCarolina,1989. M.Valtorta"OntheComplexityofBeliefNetworkRefinementandValidation."Technical ReportTR89008,DepartmentofComputerScience,UniversityofSouthCarolina,1989. M.Valtorta."TheSynthesisofInitialProbabilityAssignmentsinBeliefNetsisNP-Hard." TechnicalReportTR89007,DepartmentofComputerScience,UniversityofSouthCarolina, 1989. M.Valtorta."Knowledge-baseRefinement:ABibliography."TechnicalReportTR89006, DepartmentofComputerScience,UniversityofSouthCarolina,1989. M.Valtorta."SomeResultsonKnowledgeBaseRefinementwithanOracle."Technical ReportTR89005,DepartmentofComputerScience,UniversityofSouthCarolina,1989. M.Valtorta."SomeResultsontheComplexityofKnowledgeBaseRefinement."Technical ReportTR89004,DepartmentofComputerScience,UniversityofSouthCarolina,1989. Wagner,R.A.,C.Wright,andM.Valtorta."DetectionofUnintentionalConnectionsinVLSI Circuits."TechnicalReportCS-1984-19,DepartmentofComputerScience,DukeUniversity, 1984.

"TheLinearPlacementProblem."In:"CourseProjectsinVLSIAlgorithmics,"Technical ReportCS-1982-17,DepartmentofComputerScience,DukeUniversity,1982. 9.ResearchProposals Funded Title:"ResourceAllocationinDynamicUncertainDomains." Agency:U.S.DepartmentofDefense,DARPA. Role:PIwithJuanVargasandJoseVidal;MichaelHuhnsisProjectDirector. Amount:$779,000. Period:May15,1999toMay15,2002. Title:"DynamicDecisionSupportforCommand,Control,andCommunicationintheContext ofTacticalDefense." Agency:U.S.DepartmentofDefense. Role:Co-PrincipalInvestigator,withJohnR.Rose,SureshSingh,andAbhijitSengupta. Amount:$410,399. Period:June1,1997toJune29,2000(sic!). Note:ThisprojectispartofalargerproposalsubmittedbytheCollegeofScienceandMathematicsonFebruary17,1997. Title:"SurvivableandReconfigurableOptical/WirelessTacticalNetworks." Agency:U.S.DepartmentofDefense. Role:Co-PrincipalInvestigator,withJohnR.Rose,SureshSingh,andAbhijitSengupta. Note:Thisapproximately$400,000projectispartofalargerproposalsubmittedbytheCollegeofScienceandMathematicsonFebruary17,1997. Period:June1,1997toJune29,2000(sic!). Title:"ExpertSystemforAgriculturalLoans:CollaborationwithS.C.StateUniversity." Agency:U.S.DepartmentofAgriculture. Period:May16,1994-August31,1997. Amount:$16,289(subcontracttoUSC) Role:PrincipalInvestigator. Note:Totalamountforthe36-monthproject,entitled"AnalysisofAgriculturalLoanDefaults: DevelopmentofCredit/LoanAnalysisModels,"andtolast36months,is$296,301.Ano-cost extensionto48monthswasnegotiatedin1996. "TeachingEvaluation,Critiquing,andCurriculumChange."FundedundertheLillyTeaching FellowsProgram.IwasaJuniorTeachingFellowin1993-94intheproposedproject,whichis jointwithCarolineEastman,whoisaSeniorTeachingFellowintheproposedproject.The totalmonetaryvalueoftheawardisestimatedto$8,500. "AStudyoftheComplexityofAbstractioninQualitativeDiagnosis,"supportedbyCISESpA, January1,1991-December31,1993(withno-costextensiontoJune30,1994),for$49,500. "StudiesinSentencingInformationRetrievalandCriminalOffenseCoding,"supportedbythe SouthCarolinaLawEnforcementDivision,May4,1992-August17,1992(withMantonM. MatthewsandAbhijitSengupta),for$48,000. "DevelopmentofaComputerNetworkforQualityAssuranceandStatisticalProcessControl, Phase2,"Summer1993,GeneralElectricMedicalSystems.Mypartintheprojectwascompensatedwithapproximately$4,000.Totalfundingfortheprojectisapproximately$109,000. JuanE.Vargasisprincipalinvestigator. Submitted Title:"DistributedNormativeReasoningforMobileSensorNetworks." Agency:DepartmentofEnergy(OfficeofScienceandTechnology). Role:Co-PrincipalInvestigator(withDrs.JohnRose,MichaelHuhns,andToshiroKubota). Amount:Notavailable. Note:Pre-proposalsubmittedonJanuary2,2001.Wewerenotinvitedtosubmitafullpro-posal. Title:"NormativeDecisionAnalysisResearchIncentiveProposal." Agency:USC. Role:Co-PrincipalInvestigator(withDrs.JohnRoseandJuanVargas). Amount:Notavailable. Note:ProposalsubmittedonNovember29,2000;underreview. Title:"ITR/SI:DistributedNormativeReasoningforInteractionNetworks." Agency:NationalScienceFoundation. Role:Co-PrincipalInvestigator(withDrs.JohnRose,MichaelHuhns,andLarryStephens). Amount:Notavailable. Note:SubmittedonOctober26,2000;underreview. Title:"NormativeSupportSystemforForestManagement." Agency:InternationalPaperFoundation. Role:SoleInvestigator. Amount:Notavailable. Note:SubmittedonOctober26,2000;rejected. Title:"AnAgent-OrientedNormativeReasoningSystemforRoverPlanningandPlanExecu-tion." Agency:NASA(EpscorProgram). Role:Co-PrincipalInvestigator(withDrs.JohnRose,MichaelHuhns,andToshiroKubota). Amount:Notavailable. Note:SubmittedonSeptember22,2000.Notfunded. Title:"AnInterventionModelofNutritionandEnvironmentontheHealthoftheRural Elderly." Agency:U.S.DepartmentofAgriculture(CapacityBuildingGrantsProgram;Jointproject withS.C.StateUniversity). Role:Co-PrincipalInvestigator(withDr.KailashMathuratS.C.StateUniversity). AmountRequested:$299,250($185,567forUSC)forthreeyears. Note:SubmittedonJanuary23,1998;rejected. Title:"1998SummerInstituteforTeachersofComputerScienceAPCourses." Agency:SouthCarolinaDepartmentofEducation. Role:Co-PrincipalInvestigator(withCarterBays). Note:Rejected. Title:"AnInterventionModelofNutritionandEnvironmentontheHealthoftheRural Elderly." Agency:U.S.DepartmentofAgriculture(FundforRuralAmericaProgram). Role:Co-PrincipalInvestigator(withDrs.KailashMathuratS.C.StateandDr.JayantaK. GhoshatVorheesCollege). AmountRequested:$453,910($146,089forUSC)forthreeyears. Note:SubmittedonApril21,1997;rejected. Title:"1997SummerInstituteforTeachersofComputerScienceAPCourses." Agency:SouthCarolinaDepartmentofEducation. Role:Co-PrincipalInvestigator(withCarterBays). Note:Rejected. Title:"CombiningBayesianNetworksandCase-BasedReasoningforDiagnosisofCarcino-mas." Agency:CenterforCancerTreatmentandResearch. Role:OtherInvestigator. Note:SubmittedinNovember1994,withJuanE.Vargas.WhileIparticipatedinthepreparationofthissubmissionandwaslistedasaninvestigator,Irequestednofunding.Thisproposal wasnotaccepted. Title:Computer-AssistedScientificandLegalDecisionMakingintheCoastalDevelopment Process. Agency:NSF/EPSCoR. Role:Co-PrincipalInvestigator("TargetFaculty") Note:Otherfacultyinvolvedinthispre-proposal,submittedinJune1994,wereJamesC. EvansandDavidS.Wethey(as"mentorfaculty");TerryHuntsberger,MantonMatthews,and StephenA.Spitz(as"targetfaculty").Thisproposalwasnotaccepted. "AHighPerformanceFrameworkforATD/RbasedonWaveletsandTopologicalMatching." Submitted(inJanuary1993)toDARPA.Iamaseniorresearcherintheproposedproject. TheprincipalinvestigatorsareBjornJawerthandTerryHuntsberger.Thisproposalwasnot accepted. Isubmitted(inthefallof1990)aproposalforasmallprojectinvolvingonegraduatestudent toalargeSouthCarolinamanufacturingcompany.Thecompany,whichhadbeenencouraging atfirst,eventuallydecidednottosupportuniversityworkinthefieldofdiagnosticexpertsystems.Somedetailsofthissubmissionareconfidential,butIamavailabletodiscussitfurther onapersonalbasis. IsubmittedaproposalabstractwithS.D.Durham(StatisticsDepartment)toDARPAinthe springof1990,inanswertoDARPA/ISTOBroadAgencyAnnouncement90-12onKnowledge-BasedPlanningandScheduling.Wewererecommendednottosubmitafullproposal. Isubmittedapre-proposalwithS.D.Durham(StatisticsDepartment)toInternationalChip CorporationintheSpringof1990.Theproposalhadtobecutsosubstantiallyforbudgetary reasonsthatwedecidedtowithdrawit. Isubmitted(inthefallof1989)anapplicationforagrantentitled"AFormalAnalysisof KADS"totheUSCCommitteeonResearchandProductiveScholarship.Theproposalwas notaccepted. IsentadraftproposalinJune1989foracollaborativeNATOgranttoacolleagueinScotland whohadexpressedinterestincooperating.IreceivedapositivereplyfromhiminJanuary 1990,butIamnotfollowingupatthistime. 10.AcademicSeminars/Colloquia "SoftEvidentialUpdateforProbabilisticMultiagentSystems."DepartmentofComputerScienceandEngineering,UniversityofSouthCarolina,Columbia,S.C.,November9,2000. "TheRumorProbleminAgent-EncapsulatedBayesianNetworks"(informaltalk).Decision SupportSystemsGroup,DepartmentofComputerScience,AalborgUniversity,Aalbork,Denmark,May31,2000. "SoftEvidentialUpdateforMultiagentSystems"(informaltalk).DecisionSupportSystems Group,DepartmentofComputerScience,AalborgUniversity,Aalbork,Denmark,March30, 2000. "SomeTechniquesofDataMininginBiology."DecisionSupportSystemsGroup,Department ofComputerScience,AalborgUniversity,Aalbork,Denmark,March9,16,and23,2000(three lectures).(SomeofthismaterialwasusedfortalksattheBioinformaticsJournalCluborganizedbyDr.AustinHughesattheUniversityofSouthCarolina.Thefirstsuchtalkwasgiven onApril17,2000.) "LessonsfromtheAalborgTrip"(informaltalk).DecisionAnalysisGroup,Departmentof ComputerScienceandEngineering,UniversityofSouthCarolina,Columbia,S.C.,November8, 1999. "BuildingaBayesianNetworktoAssesstheRiskofMentalRetardationinInfants:theMEN-TORSystem."DecisionSupportSystemsGroup,DepartmentofComputerScience,Aalborg University,Aalbork,Denmark,October29,1999. "Agent-EncapsulatedBayesianNetworks."DecisionSupportSystemsGroup,Departmentof ComputerScience,AalborgUniversity,Aalbork,Denmark,October15,1999. "ValueofInformationinMultiagentEnvironments"(informaltalk).DecisionSupportSystems Group,DepartmentofComputerScience,AalborgUniversity,Aalbork,Denmark,October13, 1999. "OntheComplexityofBayesianNetworkSynthesisandRefinement."DecisionSupportSystemsGroup,DepartmentofComputerScience,AalborgUniversity,Aalbork,Denmark,October8,1999. Sixpresentationsasco-organizer(withDr.JohnRose)oftheDecisionAnalysisGroupjournal club(seehttp://www.cse.sc.edu/research/dag/journalclub.html)inFall1998,Spring1999, andFall1999. "AShortIntroductiontoBayesianNetworks."DepartmentofComputerScience,Universityof SouthCarolina,Columbia,S.C.,March26,1998. "AReviewofExistingInstrumentsfortheEvaluationofTeachingatUSC."TeachingBreakfast,October8,1996(withGailWagner,HarrietWilliams,andSteveWhisnant). "AnIntroductiontoBayesianNetworks."DepartmentofFamilyMedicine,Schoolof Medicine,UniversityofSouthCarolina,Columbia,November2,1995.

Onepresentation(onMarkovpropertiesofundirectedgraphsandHammersleyandClifford’s factorizationtheorem)intheseminarontheGibbssamplerorganizedbytheStatisticsDepartmentatUSC,March21,1995. "BayesianBeliefNetwork:AnIntroductionandSomeComputationalProblems,"Department ofPhysics,UniversityofSouthCarolina,Columbia,S.C.,January31,1994. "BayesianBeliefNetwork:AnIntroductionandSomeComputationalProblems,"Department ofStatistics,UniversityofSouthCarolina,Columbia,S.C.,January31,1994. "BayesianBeliefNetworkRefinementanditsComplexity,"DepartmentofComputerScience, ClemsonUniversity,Clemson,S.C.,January26,1993. "WhereDoHeuristicsComeFrom?"UnitedStatesArmyComputerScienceSchool,Ft.Gordon,GA,January21,1993. "OntheComplexityofDempster-ShaferBeliefNetworkSynthesis,"ArtificialIntelligenceSem-inar,SchoolofBusiness,UniversityofKansas,Lawrence,KS,March3,1992. Onepresentation(February1992)at"Hack’nSnack,"aninformalclubmeetingalmostevery week,andinvolvingfacultyfromtheChemicalEngineering,ComputerScience,Electricaland ComputerEngineeringDepartments,withoccasionalparticipationbymembersofother departmentsandindustryrepresentatives. "GenerationofHeuristicsforA*:ASurvey."CoastalCarolinaCollege,Conway,S.C.,October 25,1991. "GenerationofHeuristicsforA*:ASurvey."DepartmentofComputerScience,Universityof SouthCarolina,Columbia,S.C.,September26,1991. "RefinementofReducedRuleBases."DepartmentofElectricalandComputerEngineering, UniversityofSouthCarolina,Columbia,S.C.,April11,1991. Asaparticipantintheirvisitingprofessorprogram,IvisitedtheUnitedStatesArmyComputerScienceSchoolonJanuary29,1991.Itouredtheschoolandgavetwotalks:"Rule-BasedRefinementanditsComplexity,"and"BeliefNetworkRefinementanditsComplexity." "OntheComplexityofBeliefNetworkRefinement,"DepartmentofMathematicsandComputerScience,theCitadel,Charleston,S.C.,November1,1989. "OntheComplexityofBeliefNetworkRefinementandValidation,"CenterforMachineIntelligence,UniversityofSouthCarolina,Columbia,S.C.,October5,1989. SeveralpresentationswithintheJointResearchSeminaronEvidentialReasoningandInference Theory.(Seebelow.) "OntheComputationalComplexityofHeuristicEstimatesfortheA*Algorithm.(Or:Where DoHeuristicsComeFrom?)"JointEECE555/CSCI580session,December7,1989. "ResearchPolicyoftheEuropeanCommissionandEuropeanUnity."InternationalLawAssociationMeeting,USCLawSchool,Columbia,S.C.,October17,1989. Iwasco-organizeroftheinterdepartmentalresearchseminaronevidentialreasoningand inferencetheory(1989-1990,withSteveDurham,DepartmentofStatistics). IpresentedacolloquiumbasedonmydissertationatnineDepartmentsofComputerSciencein U.S.Universities. Iwasco-organizerofasemester-longseminaronextensionstotheGraduateCourseAdviser ExpertSystemattheComputerScienceDepartmentofDukeUniversityintheSpring1984. 11.IndustrialSeminars/Colloquia "OntheComplexityofBeliefNetworkRefinement:BayesianNetworks,"CISE,Milan,Italy, June1990. IgaveaninvitedpresentationatthelocaleventprecedingtheFirstTranscontinentalAISymposiumorganizedbyTexasInstrumentsin1986. IgaveapresentationontheGraduateCourseAdviserExpertSystematComputerCorporationofAmericainCambridge,MA,in1985. Igavealongpresentationduringaone-dayseminaronexpertsystemsatBecton-Dickinsonin ResearchTrianglePark,NC,in1984. 12.Reviewer/RefereeActivity IactedasreviewerforNSERC(1991)NSF(1990and1986),andtheIncentiveProgrammefor FundamentalResearchinArtificialIntelligenceoftheKingdomofBelgium(1987). IhaverefereedpapersforAICommunications,AnnalsofMathematicsandArtificialIntelligence,AppliedIntelligence,ArtificialIntelligence,IEEEExpert(nowIEEEIntelligentSystems),IEEETransactionsonFuzzySystems,IEEETransactionsonSystems,Man,andCybernetics,InternationalJournalofApproximateReasoning,InternationalJournalofCooperative InformationSystems,InternationalJournalofExpertSystems,InternationalJournalofUncertainty,Fuzziness,andKnowledge-BasedSystems,JournalofAutomatedReasoning,Journalof ArtificialIntelligenceResearch,JournalofExperimentalandTheoreticalArtificialIntelligence, JournalofLogicProgramming,andMachineLearning. Ihaveservedasrefereeforseveralconferences.RecentonesincludetheTwelfth,Thirteenth, Fourteenth,andFifteenthAnnualConferencesonUncertaintyinArtificialIntelligence (UAI-96,UAI-97,UAI-98,andUAI-99),the13thand16thInternationalJointConferenceson ArtificialIntelligence(IJCAI-93andIJCAI-99),the2000SymposiumonAbstraction,Reformu-lation,andApproximation(SARA-00),theFifthConferenceinSymbolicandQuantitative ApproachestoReasoningwithUncertainty(ECSQUARU-99),theSecond,Third,Fourth, Fifth,andTenthInternationalConferencesonIndustrialandEngineeringApplicationsofArtificialIntelligenceandExpertSystems(IEA/AIE-89IEA/AIE-90,IEA/AIE-91,IEA/AIE-92, IEA/AIE-97),the34rdACMSoutheasternConference(1996),the11thand8thEuropean ConferencesonArtificialIntelligence(ECAI-92andECAI-88),the10thSPIEApplicationsof AIConference,the1992SymposiumonAppliedComputing(SAC’92),theThirdInternational WorkshoponArtificialIntelligenceandStatistics(AIS-91),the23rdSoutheasternSymposium onSystemTheory(SSST-91),theFourthInternationalSymposiumonMethodologiesforIntel-ligentSystems(ISMIS-89).

IhavereviewedtextbooksandmonographsforW.H.Freeman,Benjamin-Cummings,Morgan-Kaufmann,Wiley(U.K.),Wiley(U.S.A.),andWest,including: aproposedtextbookonBayesiannetworksanddecisionnetworksforSpringer(Fall20003) aproposedtextbookonPatternRecognitionforWiley(Fall19954) aproposedtextbookonArtificialIntelligenceforW.H.FreemanandMorganKaufmann (February1991andFall19915); sevenchaptersofthesecondeditionofanArtificialIntelligencetextbookforBenjamin/Cummings,September19916andplansoftheproposedthirdedition(September1996) aproposedtextbookonKnowledge-BasedSystemsforMorganKaufmann(May19917); aproposedtextbookonLogicProgrammingforMorganKaufmann(September1990); theoutlineofatextbookontheconstructionofknowledge-basedsystemsforWiley(U.K.) (January1990); aproposedtextbook/monographonAutomaticTheoremProvingTechniquesforWiley(U.K.) (December1990-January1990); aproposedtextbookonArtificialIntelligenceforWiley(U.S.A)(November1989); aproposedtextbookonKnowledgeElicitationTechniquesforWiley(U.K)(June1989); aproposedtextbookonKnowledge-BasedSystemsforWestpublishingcompany(December 1988-January1989); theoutlineofaproposedmonographonsearchinArtificialIntelligenceforWiley(U.K.) (November1988). IactedasexternalrapporteurforaPhDdissertationdefendedinDecember1988atthe DepartmentofComputerScienceoftheUniversitedeParis-Sud. IreviewedReadingsinMedicalArtificialIntelligence,byW.K.ClanceyandE.Shortliffe (Addison-Wesley,1984),forACMSIGART(April1985,Number92,pp.13-14). 13.InvitedTalksatProfessionalMeetings Igaveaone-hourinvitedtalkattheSouthernRegionalCommitteeandAmericanStatistical Association1997SummerConferenceinGatlinburg,TN,onJune20,1997.Thetalkwasentitled"BayesianNetworks:AnIntroductionandSomeComputationalProblems." Iwasaninvitedpanelistatthe1995ACMComputerScienceConference,Nashville,TN,february28-March2,1995.Mycontributionwasentitled"IndustrialandAcademicMisperceptions ofAI."Areportonthepanelappearedasfollows.Manaris,Bill,RobertAiken,CrisKoutsougeras,ToshinoriMunakata,andM.Valtorta."ReportontheACMCSC-95Panel:’Artifi-cialIntelligence:FinallyintheMainstream?’"ACMSIGARTBulletin,6,3(July1995),7-11. Iwasinvitedtogiveapresentationatthe1989AnnualMeetingoftheSouthCarolina AcademyofScience(20-21April,1989).Titleofthepresentationis"EvidentialReasoningfor Knowledge-BasedSystems." 3ItestedthismanuscriptinCSCE590,Fall2000. 4ItestedthismanuscriptinCSCI780,Spring1995. 5Themanuscriptwaspassedfromonepublishertotheother.ItestedthemanuscriptinCSCI785,Fall1991. 6TwochaptersweretestedinCSCI785,Fall1991. 7ItestedthismanuscriptinCSCI786,Spring1991.

14.ContributedTalksatProfessionalMeetings8 McDermott,Suzanne,SubramaniMani,andM.Valtorta."ABayesianModelforthePredictionofMentalRetardationinNewborns."AbstractandPosterPresentation(givenbyS. McDermott)forPrevention-97:Science,TechnologyandPractice,TheFourteenthAnnual NationalPreventiveMedicineMeeting,Atlanta,GA,March20--23,1997. M.Valtorta."OntheDetectionofConflictsinDiagnosticBayesianNetworksUsingAbstraction."SymposiumonAbstraction,Reformulation,andApproximation(SARA-95),Ville d’Esterel,Canada,August16-18,1995. M.Valtorta."ANewResultontheComplexityofHeuristicEstimatesfortheA*Algorithm." The1994WorkshoponTheoryReformulationandAbstraction,JacksonHole,Wyoming,May 22-24,1994. M.IshaqZahidandIsubmittedabstractstothe22ndand23rdSoutheasternInternational ConferenceonCombinatorics,GraphTheory,andComputing(BocaRaton,FL,February3-7, 1992,andJanuary28,1993).Onthebasisoftheabstracts,entitledrespectively"Warnsdorff’s ToursoftheKnightona3xnboard,"and"Tie-BreakingRulesfor4xnWarnsdorff’s Tours,"wewereinvitedtogivetalks,butdidnot. "OntheRefinementandValidationofBeliefNets--PartI:Dempster-ShaferNetworks,"and "OntheRefinementandValidationofBeliefNets--PartII;BayesianNetworks."Impromptu talksgivenattheSecondAnnualWorkshoponComputationalLearningTheory(COLT-89), SantaCruz,CA,July31-August2,1989. 15.OtherProfessionalMeetings Memberoftheprogramcommittee,SeventeenthAnnualConferenceonUncertaintyinArtificialIntelligence(UAI-01),August2-5,2001,Seattle,Washington. Memberoftheprogramcommittee,FifteenthAnnualConferenceonUncertaintyinArtificial Intelligence(UAI-99),July30-August1,Stockholm,Sweden. Memberoftheorganizingcommittee,AAAISpringSymposium"PredictiveToxicologyof Chemicals:ExperiencesandImpactofAITools,"StanfordUniversity,March22-24,1999). Memberoftheprogramcommittee,FourteenthAnnualConferenceonUncertaintyinArtificial Intelligence(UAI-98),July24-26,Madison,Wisconsin. Memberoftheprogramcommittee,ThirteenthAnnualConferenceonUncertaintyinArtificial Intelligence(UAI-97),August1-3,Providence,RhodeIsland. Memberoftheprogramcommitteeandtutorialchair,TenthInternationalConferenceon IndustrialandEngineeringApplicationsofArtificialIntelligenceandExpertSystems (IEA/AIE-97),Atlanta,Georgia,onJune10-13. 8OnlypresentationsnotlistedinSection6arementionedhere.

Memberoftheprogramcommittee,TwelfthAnnualConferenceonUncertaintyinArtificial Intelligence(UAI-96),August1-3,Portland,Oregon. Memberoftheprogramcommittee,1992and1993ACMSymposiaonAppliedComputing (SAC’92andSAC’93),March1992,KansasCity,Missouri,andFebruary1993,Indianapolis, Indiana. Memberofthetechnicalprogramcommittee,23rdSoutheasternSymposiumonSystemsTheory(SSST-91),March1991,Columbia,SouthCarolina. Discussantofpaper"AnAnalysisofTwoProbabilisticModelInductionTechniques,"byS.L. CrawfordandRobertM.Fung,attheThirdInternationalWorkshoponArtificialIntelligence andStatistics(AIS-91),Ft.Lauderdale,FL,January2-5,1991(insubstitutionofX.C.Ling, whocouldnotcometotheworkshop). Localchairandmemberoftheprogramcommittee,ThirdInternationalConferenceonIndustrialandEngineeringApplicationsofArtificialIntelligenceandExpertSystems(IEA/AIE-90), July1990,Charleston,SouthCarolina. Memberoftheprogramcommittee,SixthAnnualUSC-CSSymposium:IntelligentSystems, March1990,Columbia,SouthCarolina. Co-organizerandco-chairofthespecialsession"ESPRIT--InternationalJointCollaborative AI,"10thInternationalJointConferenceonArtificialIntelligence(IJCAI-87),August1987, Milan,Italy. 16.ProfessionalSocieties AmericanAssociationforArtificialIntelligence(AAAI),ACM,ACMSpecialInterestGroupin ArtificialIntelligence(SIGART),ACMSpecialInterestGroupinAppliedComputing (SIGAPP),InstituteofElectricalandElectronicsEngineers(IEEE),IEEEComputerSociety, SouthCarolinaAcademyofScience,SigmaXi. 17.OtherProfessionalActivities AssociateEditor,InternationalJournalofApproximateReasoning,sinceJanuary1993. MemberoftheEditorialBoard,AppliedIntelligence,sinceJanuary1993. IwrotealettertotheeditorthatappearedinAIMagazine,16,3(Fall1995),p.4. IwrotealettertotheeditorthatappearedintheACMSIGARTBulletin,2,4(August1991), 4-5,as"Responseto’ExplicitSolutionstotheN-QueensProblemforallN.’" IcontributedanswerstoasurveyonteachingArtificialIntelligenceorganizedbyIEEEExpert inNovember1991.AreportonthesurveyappearedintheApril1992(vol.7,no.2)issue. 18.SupervisedGraduateStudentThesesandDissertations Young-GyunKimdefendedhisPh.D.dissertationsuccessfullyonDecember7,2000.The titleofhisdissertationis"Time-CriticalDecisionMakingwithCommunicatingInfluenceDia-grams." JayantaK.GhoshdefendedhisM.S.(P)thesissuccessfullyonNovember8,1999.Thetitle ofhisthesisis"AProbabilisticModelofHealthandNutritionofElderlyinSouthCarolina." ChuongDucHuyndefendedhisM.S.(P)thesissuccessfullyonMarch19,1999.Thetitleof histhesisis"ImplementationoftheValuation-BasedSystemforBayesianDecisionAnalysis." MarkBloemekedefendedhisPh.D.dissertationsuccessfullyonAugust7,1998,turnedin thefinalcopytotheGraduateSchoolonOctober30,1998,andreceivedhisdegreeonDecember14,1998.Thetitleofhisdissertationis:"AgentEncapsulatedBayesianNetworks." AshishKuthialadefendedhisM.S.(R)thesissuccessfullyonApril10,1998.Thetitleofhis thesisis:"ObjectModelingTechniqueandObject-OrientedAnalysisTechnique:AComparison oftheAnalysisPhasesandanImplementation." Jian-RongShidefendedhisM.S.(P)thesissuccessfullyonNovember14,1997.Thetitleof histhesisis:"SensitivitytoParameterandEvidenceValuesinBayesianNetworks." BhaskaraR.MooledefendedherM.S.(R)thesissuccessfullyonSeptember24,1997.The titleofhisthesisis"ParallelConstructionofBayesianBeliefNetworks." DonnaL.ShaverdefendedherM.S.(P)thesissuccessfullyonApril14,1997.Thetitleof herthesisis:"OfficeAutomationUsingaDatabaseApplication." GopalakrishnanViswanathdefendedhisM.S.(R)thesissuccessfullyonNovember25, 1996.Thetitleofhisthesisis:"ASurveyandComparisonofAlgorithmsfortheCompilation ofBayesianNetworks." LeszekPiatkiewiczdefendedhisM.S.(R)thesissuccessfullyonMarch20,1996.Thetitleof histhesisis:"OntheConstructionofaBayesianNetworkforAgriculturalLoanAssessment." RaghuBabuKorrapatidefendedhisM.S.(R)thesissuccessfullyonNovember10,1995. Thetitleofhisthesisis:"Model-BasedDiagnosisforContinuousSystemsUsingCLP(R)." MingWudefendedhisM.S.(P)thesissuccessfullyonJanuary20,1995.Thetitleofhisthesisis:"DesignandImplementationofaBayesianBeliefNetworkwithGraphicalUserInterface." Young-GyunKimdefendedhisM.S.(R)thesissuccessfullyonNovember2,1994.Thetitle ofhisthesisis:"DesignandConstructionofaNewStrawModelinBayesianNetworks." SubramaniManidefendedhisM.S.(P)thesissuccessfullyonAugust26,1994.Thetitleof histhesisis:"MENTOR:ABayesianModelforPredictionandInterventioninMentalRetardation." EdmundL.MaierdefendedhisM.S.(R)thesissuccessfullyonAugust7,1993.Thetitleof histhesisis:"ExpertSystemforHighSchoolStudentAdvisement." MoninderSinghdefendedhisM.S.(R)thesissuccessfullyonMay28,1993.Thetitleofthe thesisis:"ConstructionofBayesianNetworkStructuresfromData." RandyMechlingdefendedhisM.S.(R)thesissuccessfullyonJuly1,1992.Thetitleofthe thesisis"PaCCIN:AParallelConstructorofCausalIndependenceNetworks." DavidL.HiblerdefendedhisPh.D.dissertationsuccessfullyonJanuary24,1992.Thetitle ofthedissertationis"TheThoughtExperimentMethod:ANewApproachtoQualitativeReasoning." EdwardK.YudefendedhisM.S.(R)thesissuccessfullyonNovember26,1991.Thetitleof thethesisis"MODIC:AProgramforModel-BasedDiagnosisthatusesConstraintLogicPro-gramming." RitaL.ChildressdefendedherM.S.(R)thesissuccessfullyonApril16,1991.Thetitleof thethesisis"VerificationofExpertSystemsWritteninOPS5." M.IshaqZahiddefendedhisM.S.(R)thesissuccessfullyonSeptember29,1990.Thetitle ofthethesisis"Warnsdorff’sToursofaKnight." DahaiZangdefendedhisM.S.(R)thesissuccessfullyonJuly6,1990.Thetitleofthethesis is"AnAnalysisofRule-StrengthRefinementAlgorithmsforExpertSystems." ShijieWangdefendedhisM.S.(R)thesissuccessfullyonNovember30,1989.Thetitleofthe thesisis"BELFUN--ABeliefFunctionExpertSystemShell." 19.Teaching:Activities,Courses Fall2000:CSCE330,CSCE590(TopicsinInformationTechnology:FoundationsofBayesian Networks). OnsabbaticalleaveduringFall1999andSpring2000. Spring1999:CSCI330. Fall1998:CSCI220andCSCI509(TopicsinComputerScience:FoundationsofBayesian Networks9). Spring1998:Knowledge-BasedSystems(CSCI786). Fall1997:IntroductiontoAlgorithmicDesignII(CSCI146). Spring1997:CSCI330andCSCI886(SeminarinArtificialIntelligence:Decision-Theoretic ApproachestoIntelligentSystems10). Fall1996:CSCI220 Spring1996:CSCI750andCSCI886(SeminarinArtificialIntelligence:PACLearning11). 9CSCI509wastaughtasavoluntaryoverload. 10CSCI886wastaughtasavoluntaryoverload. 11CSCI886wastaughtasavoluntaryoverload.

Fall1995:CSCI780. Spring1995:CSCI146andCSCI886(SeminarinArtificialIntelligence:BayesianBelief Networks12). Fall1994:CSCI145. Spring1994:CSCI146andCSCI886(SeminarinArtificialIntelligence:Foundationsof BayesianBeliefNetworks13). Fall1993:CSCI785. Spring1993:CSCI330andCSCI886(SeminarinArtificialIntelligence:Model-BasedDiagno-sis). Fall1992:CSCI220,CSCI785,andCSCI886(SeminarinArtificialIntelligence:Constructionof BayesianBeliefNetworks14). Spring1992:Knowledge-BasedSystems(CSCI786)andDataStructuresandAlgorithms (CSCI220). Fall1991:HeuristicTechniquesinArtificialIntelligence(CSCI785)andAlgorithmsandData Structures(CSCI720). Spring1991:Knowledge-BasedSystems(CSCI786)andProgrammingLanguageStructures (CSCI330). Fall1990:HeuristicTechniquesinArtificialIntelligence(CSCI785)andProgrammingLanguageStructures(CSCI330). Spring1990:Knowledge-BasedSystems(CSCI786andCSCI786J).ThiswasaTVcourse. Fall1989:ArtificialIntelligence(CSCI580). Spring1989:Knowledge-BasedSystems(CSCI786). Fall1988:HeuristicTechniquesinArtificialIntelligence(CSCI785). Summer1984:DataStructures(CPS152,DukeUniversity). Summer1981:IntroductiontoProgramming(CPS51,DukeUniversity). 20.ServiceActivities DirectorofUndergraduateStudies,DepartmentofComputerScience,UniversityofSouth 12CSCI886wastaughtasavoluntaryoverload. 13CSCI886wastaughtasavoluntaryoverload. 14CSCI886wastaughtasavoluntaryoverload.

Carolina,Columbia,June1993-August1999.Thisdutywascompensated. Mentor,UniversityFirst-YearScholarsProgram(sinceFall1998). ActingFacultySenator,Spring1999(sinceFebruary). AdvisoroftheYear,CollegeofScienceandMathematics,UniversityofSouthCarolina, Columbia,1997. Member,UniversityCommitteeonInstructionalDevelopment,UniversityofSouthCarolina (Fall1994-Spring1997);chairoftheTeachingEvaluationsubcommittee,Fall1994-Spring1996. Chairofthefollowingdepartmentalcommittees:assessment(CSABpreparation)(Fall1992), colloquium(Spring1997,Fall1996,Spring1996,Fall1995,Spring1995,Fall1994,Spring 1994,Fall1993,Spring1993,Fall1992,Spring1991,Fall1990),curriculum(Spring1995,Fall 1994,Spring1994,Fall1993,Spring1993,Fall1992),graduatestudentadmission(Spring1992, Fall1991),qualifyingexamoversight(Spring1998,Fall1997),readingroom(Spring1998,Fall 1997),undergraduate(Spring1999,Fall1998,Spring1998,Fall1997,Spring1997,Fall1996, Spring1996,Fall1995,Spring1995,Fall1994,Spring1994). Memberofthefollowingdepartmentalcommittees:colloquium(Spring1992,Fall1991),curriculum(Spring1999,Fall1999,Spring1998,Fall1998,Spring1997,Fall1996,Spring1996, Fall1995),ethics(Spring1992,Fall1991),fullprofessorvisitation(Spring1992),graduate (Spring1991,Fall1990,Spring1990,Fall1989),graduatestudentadmission(Spring1993,Fall 1992,Spring1991,Fall1990,Spring1990,Fall1989,Spring1989,Fall1988),peerreviewand tenureandpromotionofassistantprofessors(Spring1999,Fall1998,Spring1998,Fall1997, Spring1997,Fall1996,Spring1996,Fall1995,Spring1995,Fall1994),qualifyingexam(Fall 1998,Spring1999),self-study(Spring1989),symposium(Spring1990,Fall1989,Spring1989, Fall1988). ModeratorandJudge,AlphaLambdaDeltaFreshmanHonorSociety’sSeventh,Sixth,Fifth, Fourth,andThirdAnnualUSCChallengeQuizBowl,Columbia,SC,1997,1996,1995and 1994. Judge,42nd,41st,40th,39th,38thand37thAnnualRegionIIScienceFair,Columbia,SC, 1998,1997,1996,1995,1994and1993. Graduatestudentrepresentative,DukeUniversity,DepartmentofComputerScience,1984-85. GraduatestudentrepresentativeonthechairmansearchcommitteefortheDepartmentof ComputerScience,DukeUniversity,1984.